A battery of internal resistance 0.5 ohms is connected to a load of resistance 4 ohms. If the battery has an EMF of 6.0 volts, what is the terminal voltage when it is under load?
Answer #1

let
r = 0.5 ohms
R = 4 ohms
Emf = 6.0 V

current through the battery, I = Emf/Req

= Emf/(r + R)

= 6/(0.5 + 4)

= 1.33333 A

terminal voltage, V_terminal = Emf - I*r

= 6 - 1.33333*0.5

= 5.33 V <<<<<<<<-------------------Answer